### v2

Abstract
Recent Multimodal Large Language Models (MLLMs) excel in vision-language understanding but face challenges in adapting to dynamic real-world scenarios that require continuous integration of new knowledge and skills. While continual learning (CL) offers a potential solution, existing benchmarks and methods suffer from critical limitations. In this paper, we introduce MLLM-CL, a novel benchmark encompassing domain and ability continual learning, where the former focuses on independently and identically distributed (IID) evaluation across evolving mainstream domains, whereas the latter evaluates on non-IID scenarios with new model abilities. Methodologically, we propose preventing catastrophic interference through parameter isolation and an MLLM-based routing mechanism. Extensive experiments demonstrate that our approach can integrate domain-specific knowledge and functional abilities with minimal forgetting, significantly outperforming existing methods. Our benchmark and code are available at https://github.com/bjzhb666/MLLM-CL.
Figure 1: Demonstrations of MLLM-CL benchmark. It incorporates Domain Continual Learning (DCL), which adds domain-specific knowledge, and Ability Continual Learning (ACL), which improves fundamental abilities for multimodal large language models.
1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) (Liu et al., 2024a; Chen et al., 2024b) have demonstrated remarkable capabilities in vision-language understanding. These models typically undergo supervised finetuning on carefully curated multi-task datasets, whereas real-world applications require continuous adaptation to evolving user requirements and dynamic data streams with shifting domain distributions. To incorporate new knowledge and skills, full retraining of large models is costly in both time and computing resources; besides, straightforward finetuning on novel tasks often results in catastrophic forgetting (McCloskey & Cohen, 1989; Zhai et al., 2023). Therefore, for deployment in ever-changing environments, there is an urgent need to develop MLLMs capable of continually consolidating new skills while maintaining performance on prior tasks. Recently, a few studies (Chen et al., 2024a; Zeng et al., 2024; Cao et al., 2024; Guo et al., 2025a; He et al., 2023) have explored continual learning (CL) of MLLMs. However, current works still have key limitations in both benchmarks and methodologies, preventing them from effectively exploring CL in MLLMs.
Firstly, there is a lack of well-established benchmarks. Chen et al. (2024a) proposed the first continual instruction tuning benchmark for MLLMs comprising several downstream datasets, while some of them have already been learned during the early supervised finetuning (SFT) phase of MLLM. Huai et al. (2025) divided VQAv2 (Goyal et al., 2017) into several tasks and conducted continual instruction tuning directly from the LLaVA (Liu et al., 2023) base model. However, in real-world applications, continually learning subsets of a specific dataset is impractical, and it is unlikely to finetune an MLLM on downstream tasks without any SFT on general multimodal data. Moreover, those benchmarks only consider independently and identically distributed (IID) evaluation (the training and test sets are split from the same dataset), while the model would encounter non-IID inputs in practice.
Secondly, existing methods have notable limitations: (1) Some approaches share the same set of parameters for different tasks (Chen et al., 2024a; Huang et al., 2024). This might be suitable for a conventional class-incremental learning scenario where different tasks often belong to the same dataset. However, MLLMs often encounter inputs from various domains, and the inherent task conflicts (Wei et al., 2025; Yang et al., 2024) would lead to loss of plasticity during continual learning, particularly when handling heterogeneous modalities across divergent domains. (2) Parameter isolation methods have to determine which task-specific parameters to apply for a given input during inference. This selection is usually driven by simple hand-crafted similarity metrics (Zeng et al., 2024; Guo et al., 2025a), which can be unreliable when confronted with complex multimodal data, consequently undermining overall performance.
In this paper, we establish a novel benchmark MLLM-CL, which includes two practical settings, i.e., domain continual learning (DCL) and ability continual learning (ACL), as shown in Fig. 1. Specifically, DCL aims to equip the model with domain-specific knowledge continually by learning and evaluating on several mainstream domains (remote sensing, medical, autonomous driving, science, and finance), where the training and test sets are IID. Differently, ACL focuses on incorporating fundamental abilities (OCR, math & logic, visual perception, and GUI agent), which are evaluated on non-IID test sets. Together, these two settings provide a comprehensive and realistic evaluation for continual learning of MLLMs.
Further, we design a novel method to build an efficient, lifelong-evolving MLLM. For plasticity preservation, we employ domain or ability-specific Low-Rank Adaptation (LoRA) modules (Hu et al., 2021) that maintain parameter isolation across sequentially arriving tasks, enabling comprehensive acquisition of new knowledge while preventing catastrophic interference through explicit architectural decoupling. Concurrently, to enhance parameter selection accuracy in complex multimodal scenarios, we devise a multimodal routing mechanism that leverages the model’s intrinsic multimodal understanding capabilities to automatically align input patterns with optimal task parameters. This strategy effectively transforms the MLLM’s knowledge into an explicit expert selector.
In summary, our main contributions are as follows:
  • We establish a novel benchmark for CL of MLLMs, with practical domain and ability continual learning settings, focusing on both IID and non-IID evaluation.
  • We propose a simple yet effective method with domain or ability-specific low-rank adaptation and large multimodal model-based parameter selection.
  • Experiments show that our method achieves impressive results on both domain and ability settings of the MLLM-CL benchmark, significantly outperforming existing approaches.
2 Related Work
Continual Learning.
Researchers have developed primarily four main strategies for continual learning: rehearsal-based methods (Lavda et al., 2018; Buzzega et al., 2020), regularization-based methods (Kirkpatrick et al., 2017; Li & Hoiem, 2017), structure-based methods (Mallya et al., 2018; Douillard et al., 2022), and prompt-based methods (Wang et al., 2022; Smith et al., 2023). CL in large language models has recently gained much attention (Wu et al., 2024; Shi et al., 2024a). According to the training stages, we can divide them into continual pre-training (Jang et al., 2022; Cossu et al., 2024), continual instruction tuning (Razdaibiedina et al., 2023; Zan et al., 2022; Yin et al., 2022; Wang et al., 2023a), and continual alignment (Zhang et al., 2024a; Suhr & Artzi, 2024). However, few studies focus on continual learning of MLLMs (Chen et al., 2024a; Zeng et al., 2024; Cao et al., 2024; Guo et al., 2025a; c). These prior attempts establish benchmarks with a simple dataset incremental setting where training and test sets are distributed independently and identically. Some works focus on conducting continuous instruction tuning directly from the model after the pretraining process (Huai et al., 2025; He et al., 2023). While these efforts have advanced the development of continual learning for MLLMs to some extent, they exhibit an apparent gap with the real-world production environment. Therefore, our work fills this gap and proposes a comprehensive and practical benchmark, including adding domain-specific knowledge and general abilities for CL of MLLM.
Multimodal Large Language Models.
Recent advances in MLLMs have demonstrated remarkable capabilities in multimodal understanding, open-ended generation, and instruction following across modalities. Early efforts, such as LLaVA (Liu et al., 2023; 2024a) and Qwen-VL (Bai et al., 2023), use image encoders (Radford et al., 2021) and projectors to transfer multimodal inputs into language embedding space. Recent advances (OpenAI, 2024; Li et al., 2024a; Bai et al., 2025; Fu et al., 2025) expand the ability of MLLM into more modalities, such as video and audio. With the rapid growth of MLLMs, the costs associated with training from scratch have increased dramatically (Li et al., 2024a; Tong et al., 2024; Bai et al., 2025; Chen et al., 2024c). Therefore, adapting MLLMs to dynamic environments by retraining them from scratch becomes expensive and inefficient, creating an imperative demand for continual learning of MLLMs.
3 MLLM-CL Benchmark
In this section, we provide the problem formulation and introduce the continual learning benchmark MLLM-CL. Based on the general ability and domain-specific knowledge updated in the instruction tuning stage, we divide our benchmark into domain continual learning and ability continual learning, respectively. In domain continual learning, we desire the model to learn knowledge continually, and the training sets and the test sets are IID. While in ability continual learning, we desire the model to enhance different abilities from the training data and generalize to non-IID test sets.
Problem Statement. Continual learning in MLLMs involves sequentially learning a series of multimodal tasks. Let $\mathcal{X}^{\text{img}}$ and $\mathcal{X}^{\text{ins}}$ denote the image and instruction spaces, respectively, and $\mathcal{Y}$ represent the label space for answers composed of $L$ tokens. Given a sequence of datasets $\mathcal{D}_{1},\ldots,\mathcal{D}_{T}$, where each $\mathcal{D}_{t}=\{(x^{\text{img}}_{t,i},x^{\text{ins}}_{t,i},y_{t,i})\}_{i=1}^{N_{t}}$ contains $N_{t}$ image-instruction-answer triplets drawn IID from the task-specific distribution $\mathcal{P}_{t}=\mathcal{X}_{t}^{\text{img}}\times\mathcal{X}_{t}^{\text{ins}}\times\mathcal{Y}_{t}$. Our goal is to continually update a multimodal model on observed data while retaining knowledge from previous tasks. Denote the model by $f$ with parameters $\theta_{t}$ at stage $t$, the training objective of MLLM is to predict the next token in an autoregressive way:
$\mathcal{L}_{\mathrm{MLLM}}(\theta_{t})=-\sum_{i=1}^{N_{t}}\sum_{l=1}^{L}\log p_{\theta_{t}}(y_{t,i}^{l}|x^{\text{img}}_{t,i},x^{\text{ins}}_{t,i},y_{t,i}^{<l}).$ (1)
At inference time, given an image-instruction pair $(x^{\text{img}},x^{\text{ins}})$ drawn from all learned task distributions $\{\mathcal{P}_{j}\}_{j=1}^{t}$, the model generates tokens autoregressively, i.e., the $l$-th output token is $\hat{y}^{l}=\mathop{\arg\max}\limits_{v\in\mathcal{V}}~p_{\theta}(v|x^{\text{img}},x^{\text{text}},\hat{y}^{<l})$. The above describes a typical IID scenario (e.g., domain-specific evaluation) where training and test data belong to $\{\mathcal{P}_{j}\}_{j=1}^{t}$. In practice, the model can encounter various out-of-distribution inputs $\{\mathcal{P}_{j,\text{non-iid}}\}_{j=1}^{t}\neq\{\mathcal{P}_{j}\}_{j=1}^{t}$ (e.g., ability evaluation where the input images and instruction style can be diverse), and the model is supposed to handle such a non-IID scenario.
Table 1: Statistics of the training datasets and test datasets for domain continual learning and ability continual learning. In domain continual learning, "RS" stands for remote sensing, "Med" is medical, "AD" is autonomous driving, "Sci" stands for science, and "Fin" means finance. In ability continual learning, "M & L" stands for math & logic. "VP" means visual perception.
Task Train Dataset Test Dataset Train Number Test Number
Domain Continual Learning
RS RSVQA RSVQA 60k 10k
Med PathVQA PathVQA 22.8k 9.8k
AD DriveLM DriveLM 60k 10k
Sci
AI2D, SciVerse
MapQA, TQA
AI2D, SciVerse
MapQA, TQA
33.4k
(12.4k, 0.9k, 9.6k, 7.8k)
8.2k
(3.1k, 0.2k, 2.4k, 1.9k)
Fin StockQA StockQA 60k 10k
Ability Continual Learning
OCR Monkey OCRBench 128.1k 1k
M & L MathV360K, MAVIS MathVista 526.1k 1k
VP CLEVR, TallyQA CV-Bench 119.9k 0.8k
GUI Agent
ScreenQA, MultiUI
Screen2Words
MMTBench 147.3k 0.8k
Domain Continual Learning (DCL). Continually adding domain knowledge is crucial for constructing a powerful MLLM. To achieve this goal, we propose domain continual learning and choose five mainstream and common domains: remote sensing, medical, science, autonomous driving, and finance. Specifically, we choose RSVQA (Lobry et al., 2020), PathVQA (He et al., 2020), DriveLM (Sima et al., 2023), FinVis (Wang et al., 2023b), AI2D (Kembhavi et al., 2016), SciVerse (Guo et al., 2025e), MapQA (Chang et al., 2022) and TQA (Kembhavi et al., 2017). However, FinVis is a
Figure 2: The questioner-inspector data pipeline for generating StockQA instruction tuning dataset.
caption dataset in Chinese, which may result in a language gap and is not convenient for evaluation. Therefore, we regenerate the SFT and test data as multi-choice questions and yes-or-no questions using a questioner-inspector data pipeline. Fig.˜2 shows the overall data pipeline. We use two agents, a QA generator and an inspector. Considering the varying task difficulties, we use Qwen2.5-VL-72b (Bai et al., 2025) to generate multiple choice QA pairs and Qwen2.5-VL-7b to generate Y/N QA pairs. For the inspector, we use Qwen2.5-VL-7b to check the correctness of each QA pair. After initial inspection, rule-based formatting is applied to generate the final dataset, named StockQA. All experiments are conducted using the vllm (Kwon et al., 2023) engine. Appendix˜B provides detailed prompts for each agent, rules for filtering, examples, and statistics of the StockQA dataset. Tab.˜1 shows the statistics of the datasets for DCL and Fig.˜1 shows some examples. More examples are provided in the Sec.˜G.1.
Ability Continual Learning (ACL). DCL assumes that training and test data are IID. However, achieving IID between training and test sets is often challenging in real-world scenarios, which has been ignored by existing benchmarks (Chen et al., 2024a; Zeng et al., 2024; Guo et al., 2025a; Cao et al., 2024). Therefore, we consider a more challenging setting with non-IID training and test data, which we term ability continual learning. For ACL, we select four fundamental abilities for the MLLM to learn sequentially: OCR, math & logic, visual perception, and GUI agent. In terms of the SFT data, we collect the training data from LLaVA-OneVision (Li et al., 2024a), Monkey (Li et al., 2024b), ScreenQA (Hsiao et al., 2022), Screen2Words (Wang et al., 2021), MultiUI (Liu et al., 2024b), Math-LLaVA (Shi et al., 2024b), MAVIS (Zhang et al., 2024b), CLVER (Johnson et al., 2017) and TallyQA (Acharya et al., 2019) and testing data from OCRBench (Liu et al., 2024d), MathVista (Lu et al., 2024), MMTBench-GUI (Ying et al., 2024) and CV-Bench-Counting (Tong et al., 2024), respectively. Tab.˜1 presents the details of the datasets for training and testing in ACL, and Fig.˜1 provides a demonstration. Additional examples can be found in the Sec.˜G.1.
4 The Proposed Method: MR-LoRA
4.1 Training: Expert Learning without Task Conflict
Figure 3: Prompt of the MLLM-based router selector.
Figure 4: Comparison of new task performance (LLaVA-based) on both domain and ability CL.
Learning Low-Rank Expert without Task Conflict. In traditional continual learning, particularly class-incremental learning, the model for learning a new task is typically initialized with parameters from the previous task to facilitate knowledge transfer, and then various regularization constraints are incorporated to mitigate catastrophic forgetting. Therefore, a natural question arises: Is this paradigm suitable for continual learning in MLLMs? Some studies (Wei et al., 2025; Yang et al., 2024) have revealed that data interference widely exists in the training of MLLMs. We empirically investigate the task conflict problem of domain and ability continual learning by comparing the average new task performance. The results in Fig.˜4 yield the following observation: (1) Initializing with weights from prior tasks (e.g., LoRA-FT, MoELoRA (Chen et al., 2024a)) reduces model plasticity, leading to worse performance than learning each task individually with randomly initialized LoRA (i.e., scratch). (2) Regularization (e.g., O-LoRA (Wang et al., 2023a), SEFE (Chen et al., 2025)) or parameter-sharing-based methods (e.g., CL-MoE (Huai et al., 2025), HiDE (Guo et al., 2025a)) also suffer from loss of plasticity when learning new tasks. (3) The task conflict in DCL is more severe than that in ACL, which is reasonable because the domain gap in DCL (e.g., autonomous driving vs. science) is often larger than that in ACL (OCR vs. Math). Based on the above analysis, we propose initializing a fresh LoRA (Hu et al., 2021) module from scratch for each task to circumvent inter-task conflicts when learning new domains. Compared to the original parameters of the large model, LoRA introduces minimal additional parameters, enabling domain-specific adaptation via lightweight, task-exclusive adapters.
Few-shot Router Tuning. In our framework, we tune a low-rank expert for each domain or capability, and dynamically select the most appropriate expert at inference time. While existing selection strategies (Zeng et al., 2024; Guo et al., 2025a) rely on simple similarity measures, e.g., computing cosine similarity between task prototypes and sample features in the embedding space, multimodal scenarios involve more complex inputs. Therefore, we propose leveraging the MLLM’s intrinsic capability to process complex multimodal inputs by tuning an MLLM-based selection router. This router identifies the corresponding expert for each input. Specifically, for each task, we collect a few-shot set $\mathcal{M}_{t}=\{(x^{\text{img}}_{t,i},x^{\text{ins}}_{t,i})\}_{i=1}^{m}$, where $m\ll N_{t}$ (we set $m=20$ in all experiments). After each continual learning phase, the accumulated few-shot data $\{\mathcal{M}_{j}\}_{j=1}^{t}$ and expert model descriptions are transformed into structured instructions. We adopt a generative style to select the most suitable expert and tune the MLLM using a router LoRA via autoregressive loss (Liu et al., 2024a). An illustration of the router selection prompt for domain continual learning is provided in Fig.˜3.
4.2 Inference: Router Selection with MLLM
Figure 5: Overall framework of our MR-LoRA.
Framework of MR-LoRA. During inference, with expert learning and router selection, the overall framework of the proposed method is illustrated in Fig.˜5. Our MR-LoRA performs two-stage inference for a given multimodal input, consisting of a routing phase followed by a prediction phase. In the first stage, the expert selection router is performed to select a domain or ability-specific expert. Then, the selected expert is combined with the pre-trained backbone to output the final response. On the one hand, by decoupling the learning of different domains or abilities, we avoid potential distribution conflict and can learn a good expert for a given task. On the other hand, the proposed router selection strategy largely explores the advantages of MLLMs to improve the flexibility and accuracy of expert selection, ensuring promising final prediction performance during continual learning. The proposed MLLM-based routing mechanism offers notable advantages: (1) The MLLM’s strong multimodal understanding capacity ensures robust expert selection performance on complex multimodal inputs. (2) The selection router is parameter-efficient and learned with few-shot unlabeled image-question pairs, allowing on-the-fly adaptation.
5 Experiments
5.1 Experimental Setup
[…]
Evaluation Metric. We report the last accuracy, which is the accuracy of all seen tasks after learning the last task, mean finetune accuracy (MFT), mean final accuracy (MFN), mean average accuracy (MAA), and backward transfer (BWT) following standard metrics in continual learning (Guo et al., 2025a; Chen et al., 2025). The detailed calculation of each metric is shown in the Sec.˜A.3.
Table 2: Results for LLaVA-based domain continual learning in MLLM-CL benchmark. denotes the original method with replay data.
[…]
Zeroshot 32.29 28.28 15.59 35.55 62.56 34.85 - - -
Oracle 81.06 65.83 54.17 56.86 91.14 69.81 - - -
LoRA-FT (Hu et al., 2021) 69.65 41.59 25.43 40.88 87.45 64.98 53.00 61.13 -14.97
LoRA-FT (Hu et al., 2021) 76.54 50.27 43.01 43.32 89.85 66.32 60.60 64.72 -7.15
O-LoRA (Wang et al., 2023a) 74.64 44.42 30.02 41.47 87.15 65.16 55.54 62.12 -12.03
O-LoRA (Wang et al., 2023a) 76.94 41.17 34.18 39.61 83.22 60.49 55.02 60.73 -6.83
MoELoRA (Chen et al., 2024a) 77.54 41.85 27.62 40.13 86.75 64.94 54.78 61.76 -12.70
MoELoRA (Chen et al., 2024a) 77.63 49.54 39.08 41.04 89.21 66.24 59.30 64.81 -8.68
CL-MoE (Huai et al., 2025) 71.34 46.84 26.33 41.17 88.74 66.06 54.88 61.79 -13.96
CL-MoE (Huai et al., 2025) 76.58 52.31 39.65 45.64 90.21 66.65 60.88 64.95 -7.22
HiDe (Guo et al., 2025a) 74.31 48.95 33.21 38.54 81.55 60.77 55.31 60.68 -6.82
HiDe (Guo et al., 2025a) 74.80 42.29 34.03 38.01 79.22 60.83 53.67 61.81 -8.95
SEFE (Chen et al., 2025) 77.26 50.37 37.21 40.87 86.82 65.01 58.51 63.63 -8.13
SEFE (Chen et al., 2025) 78.43 52.85 46.21 47.76 89.33 66.89 62.92 66.51 -4.97
DISCO (Guo et al., 2025b) 76.03 45.20 43.79 42.33 88.95 64.43 59.26 63.35 -6.46
DISCO (Guo et al., 2025b) 77.78 46.25 50.45 49.51 89.71 65.27 62.74 64.92 -3.17
MR-LoRA (Ours) 80.87 65.32 54.12 56.71 91.12 69.64 69.63 71.06 -0.01
Table 3: Results for LLaVA-based ability continual learning in MLLM-CL benchmark.
[…]
Zeroshot 31.20 30.20 60.79 10.00 33.05 - - -
Oracle 33.60 36.50 65.10 32.50 41.93 - - -
LoRA-FT (Hu et al., 2021) 23.60 33.70 55.84 32.50 41.28 36.41 36.58 -6.49
LoRA-FT (Hu et al., 2021) 21.80 32.70 58.38 28.75 40.32 35.41 36.32 -6.55
O-LoRA (Wang et al., 2023a) 29.60 32.90 52.41 33.75 39.72 37.16 35.42 -3.41
O-LoRA (Wang et al., 2023a) 29.60 31.30 60.79 27.50 39.96 37.30 36.34 -3.55
MoELoRA (Chen et al., 2024a) 26.70 32.80 56.85 27.22 39.45 35.89 36.07 -4.75
MoELoRA (Chen et al., 2024a) 19.80 32.20 54.19 30.00 40.35 34.05 35.39 -8.41
CL-MoE (Huai et al., 2025) 19.90 32.70 53.43 30.69 40.50 34.18 35.65 -8.43
CL-MoE (Huai et al., 2025) 25.40 31.80 60.91 30.00 41.22 37.03 37.28 -5.59
HiDe (Guo et al., 2025a) 24.60 32.10 46.32 28.75 37.98 32.94 34.60 -6.72
HiDe (Guo et al., 2025a) 24.60 28.40 30.71 23.75 36.84 26.86 33.54 -13.30
SEFE (Chen et al., 2025) 26.00 33.40 57.74 33.75 40.98 37.72 36.59 -4.35
SEFE (Chen et al., 2025) 25.60 34.80 57.61 31.39 42.25 37.35 37.93 -6.53
DISCO (Guo et al., 2025b) 32.90 33.10 60.15 30.14 39.02 39.07 36.57 0.07
DISCO (Guo et al., 2025b) 34.20 35.00 61.55 27.50 40.14 39.56 37.85 -0.77
MR-LoRA (Ours) 33.70 36.20 65.10 32.50 41.89 41.88 38.86 -0.02
Table 4: Results for InternVL-based domain continual learning in MLLM-CL benchmark. denotes the original method with replay data.
[…]
Zeroshot 31.16 29.81 14.06 33.93 64.32 34.66 - - -
Oracle 81.49 66.42 54.56 54.48 91.24 69.64 - - -
LoRA-FT (Hu et al., 2021) 69.93 52.17 33.04 42.67 91.07 69.06 57.78 65.22 -14.11
LoRA-FT (Hu et al., 2021) 77.06 47.55 42.67 43.31 91.44 69.43 60.41 67.45 -11.28
MoELoRA (Chen et al., 2024a) 69.90 52.08 33.17 42.19 90.58 68.83 57.58 65.97 -14.06
MoELoRA (Chen et al., 2024a) 76.74 52.65 38.81 42.15 89.84 67.90 60.04 66.01 -9.83
HiDe (Guo et al., 2025a) 75.40 57.66 36.73 41.48 88.59 65.26 59.97 65.94 -6.60
HiDe (Guo et al., 2025a) 53.17 52.61 40.85 47.04 89.17 64.20 56.57 61.06 -9.54
DISCO (Guo et al., 2025b) 75.12 50.69 52.41 50.67 90.86 68.85 63.95 68.14 -6.12
DISCO (Guo et al., 2025b) 77.90 47.50 49.13 49.37 90.92 68.55 62.96 67.81 -6.98
MR-LoRA (Ours) 81.48 65.80 54.56 54.40 91.07 69.51 69.46 71.27 -0.06
Table 5: Results for InternVL-based ability continual learning in MLLM-CL benchmark.
[…]
Zeroshot 30.00 31.20 56.09 2.50 29.95 - - -
Oracle 32.20 33.40 67.77 33.75 41.78 - - -
LoRA-FT (Hu et al., 2021) 21.40 32.80 60.28 29.86 40.84 36.08 36.38 -6.35
LoRA-FT (Hu et al., 2021) 26.30 34.20 62.56 31.25 41.63 38.58 37.38 -4.07
O-LoRA (Wang et al., 2023a) 25.50 32.30 64.59 24.44 38.64 36.71 36.05 -2.57
O-LoRA (Wang et al., 2023a) 21.70 31.10 59.77 31.25 41.38 35.96 36.49 -7.23
MoELoRA (Chen et al., 2024a) 17.20 32.70 55.33 32.50 41.41 34.43 35.36 -9.30
MoELoRA (Chen et al., 2024a) 13.90 29.70 54.95 32.50 41.91 32.76 35.66 -12.20
HiDe (Guo et al., 2025a) 17.70 33.00 41.12 20.28 37.27 28.02 33.25 -12.33
HiDe (Guo et al., 2025a) 25.30 29.20 42.13 20.28 35.93 29.23 33.39 -8.93
DISCO (Guo et al., 2025b) 30.60 33.10 65.36 27.50 39.21 39.14 36.73 -0.10
DISCO (Guo et al., 2025b) 32.30 32.30 64.97 30.14 40.46 39.93 37.63 -0.71
MR-LoRA (Ours) 33.00 35.70 67.51 33.75 42.56 42.49 38.85 -0.09
5.2 Results and Analysis
Domain Continual Learning. As demonstrated in Tab.˜2 (LLaVA-based) and Tab.˜4 (InternVL-based), our proposed MR-LoRA method achieves state-of-the-art performance on the DCL setting, showcasing its exceptional ability to acquire new domain knowledge while preserving previously learned capabilities. The performance of MR-LoRA highlights several key advantages: (1) Approaching Oracle Performance: Our method’s final accuracy on all individual tasks nearly matches the “Oracle” performance. For instance, in Tab.˜2, the final accuracies of MR-LoRA across the five domains are almost identical to the Oracle scores. This indicates that our MLLM-based router can select the most appropriate expert module for each input sample with high precision, allowing the overall performance to approach the theoretical upper bound of a perfect selection mechanism. (2) Superiority over Existing Baselines: In contrast, other baseline methods exhibit significant performance degradation. Parameter-sharing and regularization methods like LoRA-FT and O-LoRA suffer from severe forgetting, as evidenced by their deeply negative BWT scores (e.g., -14.97 for LoRA-FT on LLaVA). This empirically confirms our hypothesis in Sec.˜4.1 regarding the severe task conflict among heterogeneous domains, where shared parameters compromise existing abilities while learning new ones. Although replay-based methods (marked with ) alleviate forgetting by rehearsing old data, their performance remains far inferior to MR-LoRA. Even more advanced baselines like DISCO and SEFE still show a significant gap compared to ours.
Figure 6: Examples demonstrating that the selected expert handles certain questions better than the original expert in DCL and ACL. MLLM-enhanced router selects the most appropriate experts.
Ability Continual Learning. The effectiveness of our proposed method in the more challenging ACL setting is demonstrated in Tabs.˜3 and 5. This setting evaluates the model’s capacity to acquire fundamental new skills and generalize to non-IID test sets. Firstly, we observe that most baselines suffer from severe catastrophic forgetting, revealing a critical weakness in existing CL approaches when faced with real-world, practical non-IID scenarios. In contrast, our MR-LoRA significantly outperforms all baseline methods and successfully improves performance across all four abilities by isolating abilities into dedicated expert modules and leveraging an intelligent MLLM-based router.
Table 6: Ablation study of LoRA rank for each expert LoRA (LLaVA, DCL, last accuracy).
Rank RS (%) Med (%) AD (%) Sci (%) Fin (%)
8 80.96 64.64 54.00 55.44 90.75
16 80.92 65.11 53.98 55.90 91.02
32 80.87 65.32 54.12 56.71 91.12
64 81.18 66.07 54.31 56.90 91.60
128 81.14 66.49 54.00 57.63 91.44
Interestingly, the results also reveal a knowledge transfer enabled by our MLLM-enhanced router. In the InternVL experiments, the final accuracy of MR-LoRA on the OCR task is 33.00%, which is higher than the 32.20% achieved by the Oracle. This suggests that the router’s flexible selection mechanism can sometimes leverage knowledge from other related experts (e.g., using the OCR capabilities in the M & L expert) to achieve a result superior to that of a single, isolated specialist. This phenomenon highlights the rationality and sophisticated decision-making capability of the MR-LoRA framework. Fig.˜6 shows the knowledge transfer phenomenon in DCL and ACL.
Rank of Expert LoRA. From the results in Tab.˜6, we find that our method performs well even at very low ranks (e.g., 8), demonstrating its parameter efficiency. This indicates that even if the number of tasks to be learned is large, our method can still achieve a good performance with only a small increase in parameters. Besides, as the expert rank increases, performance can be improved slightly because of more trainable parameters.
Router Accuracy. We ablate the number of samples for routing data and report the router selection accuracy and the last accuracy in domain and ability continual learning. The results are shown in Secs.˜5.2 and 8. In DCL, we find that our method can achieve an excellent performance (almost 100% selection accuracy) using only 20 samples to train the router, which means our method closes the gap of training each task individually. Note that the number of samples we used is much smaller than the number of training samples (60k). Besides, with more sampling data, the router selection accuracy improves and the performance of MR-LoRA slightly increases. In ACL, the performance of MR-LoRA achieves satisfactory performance when the shot of router tuning is 10. It is interesting that the router accuracy of the OCR task is around 50%, but our method can achieve a comparable, or even better performance compared with directly finetuning an OCR LoRA expert (33.60%). This means MR-LoRA uses other experts to solve the OCR task, and these experts perform well on these test samples. It is reasonable that OCR is a basic and fundamental ability that the math and GUI Agent experts are also able to extract equations and web texts from the images.
Table 7: Router accuracy under different amount of router data in domain continual learning. The left part is the router selection accuracy and the right part is task accuracy after learning the last task.
# Replay
Samples
Router Accuracy (%) Last Accuracy (%)
RS Med AD Sci Fin RS Med AD Sci Fin
100 99.96 99.16 99.98 98.44 99.99 81.04 65.61 54.16 56.77 91.13
50 99.85 98.69 99.94 98.82 100.00 81.00 65.53 54.14 56.76 91.14
30 99.62 98.89 100.00 96.90 99.86 80.92 65.53 54.17 56.59 91.08
20 99.52 97.87 99.89 98.40 99.80 80.87 65.32 54.12 56.71 91.12
10 99.93 98.24 99.93 97.75 99.40 81.04 65.40 54.16 56.63 91.01
Table 8: Router accuracy under different amount of replay data in ability continual learning.
# Replay
Samples
Router Accuracy (%) Last Accuracy (%)
OCR M&L VP GUI Agent OCR M&L VP GUI Agent
100 72.10 94.60 99.87 100.00 32.80 36.30 65.10 32.50
50 65.30 83.90 99.11 100.00 32.70 36.10 64.85 32.50
30 53.60 90.90 97.21 98.38 33.80 36.70 64.85 32.50
20 51.40 86.00 100.00 100.00 33.70 36.20 65.10 32.50
10 81.90 76.30 100.00 100.00 32.80 35.80 65.10 32.50
6 Conclusion
In this paper, we first propose MLLM-CL benchmark, a novel benchmark including domain continual learning and ability continual learning. In domain continual learning, we select five specific domains (remote sensing, medical, science, autonomous driving, and finance) and focus on IID evaluation. In ability continual learning, we consider a more practical setting where the training and test sets are non-IID. We select four common and fundamental abilities for MLLM to learning sequentially: OCR, math & logic, visual perception, and GUI agent. To solve the two settings in the MLLM-CL benchmark, we first analyze the task conflict between different tasks and then propose an MLLM enhanced router selection method MR-LoRA. Comprehensive experiments and analyses validate the necessity of our MLLM-CL benchmark and show the effectiveness and efficiency of our proposed method. We believe that our carefully designed benchmark and MR-LoRA can serve as a foundation for continual learning in multimodal large language models and will introduce an innovative and practical direction of continual learning and MLLM to the community.
Ethics statement
Our research is grounded in ethical practices, with particular attention paid to the responsible use of data. This work exclusively employs public, well-established datasets from the MLLM community, and we list all used assets’ licenses in Tab.˜12. Our use of this data is in accordance with their provided licenses and intended academic purpose.
Reproducibility statement
To facilitate the reproducibility of our research, we provide comprehensive implementation details in Appendix˜A, including training procedures and hyperparameters. We also report all the result matrices in Appendix˜C. All source code, datasets, and trained models will be publicly released upon the paper’s acceptance.
Appendix
Appendix A Implementation Details
In this section, we introduce the implementation details of MR-LoRA and the evaluation details of each task in domain continual learning and ability continual learning.
A.1 Training Details
DCL. Tab.˜9 shows the hyperparameters for training the router and expert in domain continual learning. For most configurations, we follow the default setting of LLaVA 1.5 (Liu et al., 2023). To ensure comparable training exposure across datasets of varying sizes, each task is trained for approximately 60,000 instances in DCL. For efficient fine-tuning, a rank of 32 is employed. For all the experiments, we use 8 A100 GPUs, and the training time for each task is around 1 hour.
ACL. Tab.˜10 shows the hyperparameters for ability continual learning. For ability continual learning, training time is around 20 hours to train all the tasks sequentially.
Router Training. For the router training, we train 30 epochs in domain continual learning and ability continual learning; we keep other configurations identical to the experts’ except for the learning rate. We use the codebase from MCITlib (Guo et al., 2025d) and LLaVA (Liu et al., 2023).
Table 9: Hyperparameters of MR-LoRA in domain continual learning
Expert Config Router Config
LLaVA InternVL LLaVA InternVL
optimizer AdamW AdamW
batch size $64$ $64$
lr schedule cosine decay cosine decay
lr warmup ratio $0.03$ $0.03$
LoRA rank $32$ $32$
DeepSpeed stage 22 22
base lr $1\text{\times}{10}^{-4}$ $2\text{\times}{10}^{-5}$ $1\text{\times}{10}^{-4}$
epoch for RS 11 -
epoch for Med 33 $30$
epoch for AD 11 $30$
epoch for Sci 22 $30$
epoch for Fin 11 $30$
A.2 Evaluation Details
In domain continual learning, for the financial task, all the questions are MCQ or Y/N questions; we require the prediction to exactly match the ground truth. For autonomous driving, medical, and remote sensing tasks, we consider the prediction to include the ground truth as the correct answer. This serves as the default evaluation method. For science tasks, some test samples are multiple-choice questions (MCQs), and predictions are required to exactly match the ground truth. Certain questions in MapQA (Chang et al., 2022) require the model to list places; in these cases, we compute the percentage of correct responses. Other science questions are evaluated according to the default method. In ability continual learning, we follow the default setting of the corresponding benchmarks.
A.3 Detailed Evalution Metrics
We used the integrated metrics in SEFE and MCITlib (Chen et al., 2025; Guo et al., 2025d) to evaluate the performance of each method.
  • Last accuracy is the accuracy of all seen tasks after learning the last task.
  • Mean Finetune Accuracy (MFT) measures the average accuracy achieved on each task immediately after it is learned, serving as an upper bound that reflects the model’s performance in the absence of forgetting.
  • Mean Final Accuracy (MFN) computes the average accuracy over all tasks after completing the full incremental training process, representing the model’s overall retained performance.
  • Mean Average Accuracy (MAA) calculates the mean of average accuracies on all learned tasks after each training step, offering a holistic view of performance throughout the continual learning process.
  • Backward Transfer (BWT) captures the change in accuracy for each task by comparing its final accuracy with that immediately after it was learned, quantifying the extent of forgetting.
For clarity, a conceptual illustration of the evaluation metrics is provided in  Fig.˜7.
Table 10: Hyperparameters of MR-LoRA in ability continual learning
Expert Config Router Config
LLaVA InternVL LLaVA InternVL
optimizer AdamW AdamW
batch size $128$ $128$
lr schedule cosine decay cosine decay
lr warmup ratio $0.03$ $0.03$
LoRA rank $32$ $32$
DeepSpeed stage 22 22
base lr (OCR) $5\text{\times}{10}^{-5}$ $2\text{\times}{10}^{-4}$ - -
base lr (M&L,VP,GUI) $2\text{\times}{10}^{-4}$ $2\text{\times}{10}^{-4}$ $2\text{\times}{10}^{-4}$ $1\text{\times}{10}^{-4}$
epoch for OCR 33 -
epoch for Math & Logic 11 $30$
epoch for VP 11 $30$
epoch for GUI Agent 33 $30$
Figure 7: Illustration of the evaluation metric calculations (Guo et al., 2025d).
A.4 Router prompt for MR-LoRA
We previously provided our router prompt for DCL in Fig.˜3. The prompt for ACL appears in Fig.˜8.
Figure 8: Prompt for the router in ability continual learning.
Figure 9: Prompt for the Questioner to generate MCQ question answer pairs.
Figure 10: Prompt for the Questioner to generate Y/N question answer pairs.
Figure 11: Prompt for the Inspector to check the question answer pairs.
Appendix B Details of StockQA Dataset
Overview. The StockQA dataset is a multimodal financial dataset concentrated on stock analysis. It is rewritten from the FinVis (Wang et al., 2023b) dataset.
Finvis dataset is a Chinese caption dataset generated by GPT4V (Achiam et al., 2023). All the captions are related to the stock technical indicator analysis. However, the caption form is not convenient for evaluation, and there may be a language gap between this task and other tasks. Therefore, we use a questioner-inspector data pipeline with a powerful MLLM Qwen2.5-VL (Bai et al., 2025) to rewrite the caption into MCQ and Y/N question-answer pairs and name it StockQA. When manually checking the inspector process, we find that the inspector misclassified some correct question-answer pairs. Nevertheless, it successfully identified erroneous instances, thereby contributing to the overall correctness of the final dataset.
Prompts for agents. Figs.˜10 and 9 shows the prompt we use for the Questioner to generate Y/N and MCQ question-answer pairs, respectively. Fig.˜11 is the prompt we use for the inspector.
Rules for filtering. After using an inspector agent to check the correctness and rationality, we employ the following rules to balance the choices of multiple choice questions to mitigate the position bias (Liu et al., 2024c) and format the output.
  • Format: Remove the unnecessary spaces, line breaks, and punctuation to make each question in the same format.
  • Position: Exchange the choices of multiple choice questions to ensure the right answers of the total datasets are distributed with the same probability.
Table 11: Statistics of the StockQA dataset.
Data Max Length Min Length Average Length Amount
MCQ 683 115 392.74 48k
TF 99 21 42.29 22k
Total 683 21 282.60 70k
Table 12: Existing assets grouped by license.
License Assets
$C$-BY-SA-4.0 TQA, MapQA, MathVista, AI2D
Apache-2.0 DriveLM, MathV360k, CV-Bench, CoIN
MIT Monkey, OCRbench, MAVIS
$C$-BY-4.0 CLEVR, ScreenQA, Screen2Words, MMTBench
Figure 12: Word length distribution of the StockQA dataset.
Figure 13: MCQ and Y/N examples in StockQA dataset.
Statistics of StockQA dataset. StockQA is a new VQA dataset related to multimodal stock analysis. It includes 70k question-answer pairs. of which 60k is the training set and 10k is the test set. For the training data, there are 40k MCQ and 20k Y/N QA pairs. For the test data, there are 8k MCQ and 2k QA pairs. Each choice is equally distributed after our cleaning process. Figs.˜14 and 13 shows the word cloud and examples of StockQA dataset. Tabs.˜11 and 12 shows the detailed statistics of StockQA dataset.
Dataset License. Our dataset follows the CC-BY license. This license allows reusers to distribute, remix, adapt, and build upon the material in any medium or format, so long as attribution is given to the creator. The license allows for commercial use. For other assets we used, we list the licenses below in Tab.˜12.
Figure 14: Word cloud of StockQA dataset.
Appendix C Detailed Continual Learning Results
In this section, we show the detailed inference results of all the methods (LoRA (Hu et al., 2021), LoRA (Hu et al., 2021), O-LoRA (Wang et al., 2023a), O-LoRA (Wang et al., 2023a), MoELoRA (Chen et al., 2024a), MoELoRA (Chen et al., 2024a), CL-MoE (Huai et al., 2025), CL-MoE (Huai et al., 2025), HiDe (Guo et al., 2025a), HiDe (Guo et al., 2025a), SEFE (Chen et al., 2025), SEFE (Chen et al., 2025), DISCO (Guo et al., 2025b), DISCO (Guo et al., 2025b) and MR-LoRA) during each continual learning stage, where  denotes the original method with replay data.
C.1 Baseline Results in Domain Continual Learning
Table 13: Result matrices of InternVL-based baselines in domain continual learning. denotes the original method with replay data.
LoRA-FT RS Med AD Sci Fin
RS 81.29
Med 75.71 65.92
AD 69.38 56.87 53.56
Sci 71.12 53.75 46.83 53.48
Fin 69.93 52.17 33.04 42.67 91.07
LoRA-FT RS Med AD Sci Fin
RS 81.68
Med 77.45 66.69
AD 77.24 61.32 53.81
Sci 77.89 55.43 49.13 53.53
Fin 77.06 47.55 42.67 43.31 91.44
MoELoRA RS Med AD Sci Fin
RS 81.22
Med 77.56 66.00
AD 74.56 58.74 53.62
Sci 72.62 54.77 47.65 52.75
Fin 69.90 52.08 33.17 42.19 90.58
MoELoRA RS Med AD Sci Fin
RS 80.75
Med 78.10 64.77
AD 73.24 59.54 52.90
Sci 76.82 53.64 42.11 51.24
Fin 76.74 52.65 38.81 42.15 89.84
HiDe RS Med AD Sci Fin
RS 81.24
Med 79.59 64.71
AD 78.85 58.37 41.75
Sci 78.33 58.51 39.94 49.99
Fin 75.40 57.66 36.73 41.48 88.59
HiDe RS Med AD Sci Fin
RS 73.92
Med 71.44 64.22
AD 65.74 49.83 42.92
Sci 70.64 54.79 40.47 50.75
Fin 53.17 52.61 40.85 47.04 89.17
DISCO RS Med AD Sci Fin
RS 81.42
Med 79.13 63.80
AD 78.62 60.79 53.98
Sci 77.40 52.21 53.74 54.18
Fin 75.12 50.69 52.41 50.67 90.86
DISCO RS Med AD Sci Fin
RS 81.49
Med 80.14 63.05
AD 78.87 57.42 53.77
Sci 78.67 52.80 53.56 53.52
Fin 77.90 47.50 49.13 49.37 90.92
Table 14: Result matrices of LLaVA-based baselines in domain continual learning. denotes the original method with replay data.
LoRA-FT RS Med AD Sci Fin
RS 78.32
Med 74.68 57.53
AD 68.93 47.19 52.15
Sci 75.12 45.56 38.46 49.44
Fin 69.65 41.59 25.43 40.88 87.45
LoRA-FT RS Med AD Sci Fin
RS 79.33
Med 76.45 57.58
AD 74.54 54.26 52.96
Sci 77.00 50.31 45.13 51.88
Fin 76.54 50.27 43.01 43.32 89.85
O-LoRA RS Med AD Sci Fin
RS 79.25
Med 74.05 56.52
AD 76.06 43.71 52.32
Sci 76.60 44.87 40.57 50.58
Fin 74.64 44.42 30.02 41.47 87.15
O-LoRA RS Med AD Sci Fin
RS 79.17
Med 78.21 56.65
AD 77.52 38.60 37.81
Sci 77.61 44.22 35.40 45.59
Fin 76.94 41.17 34.18 39.61 83.22
MoELoRA RS Med AD Sci Fin
RS 79.09
Med 74.78 58.73
AD 77.69 43.72 51.47
Sci 76.87 43.79 32.81 48.67
Fin 77.54 41.85 27.62 40.13 86.75
MoELoRA RS Med AD Sci Fin
RS 79.66
Med 78.44 60.50
AD 78.54 49.86 52.54
Sci 78.00 50.53 43.32 49.30
Fin 77.63 49.54 39.08 41.04 89.21
CL-MoE RS Med AD Sci Fin
RS 79.08
Med 73.48 60.56
AD 72.61 44.42 51.62
Sci 71.02 48.04 37.70 50.28
Fin 71.34 46.84 26.33 41.17 88.74
CL-MoE RS Med AD Sci Fin
RS 79.40
Med 76.32 61.10
AD 72.01 54.49 52.56
Sci 76.64 53.89 43.83 49.98
Fin 76.58 52.31 39.65 45.64 90.21
HiDe RS Med AD Sci Fin
RS 78.14
Med 74.26 58.05
AD 74.90 42.94 39.65
Sci 75.43 44.91 38.33 46.44
Fin 74.31 48.95 33.21 38.54 81.55
HiDe RS Med AD Sci Fin
RS 79.21
Med 77.79 60.88
AD 77.64 48.19 38.12
Sci 77.51 48.84 35.76 46.71
Fin 74.80 42.29 34.03 38.01 79.22
SEFE RS Med AD Sci Fin
RS 78.27
Med 76.32 58.42
AD 77.22 49.13 52.49
Sci 77.83 47.70 43.01 49.04
Fin 77.26 50.37 37.21 40.87 86.82
SEFE RS Med AD Sci Fin
RS 79.21
Med 78.39 60.93
AD 79.00 57.68 53.11
Sci 78.76 51.39 47.99 51.87
Fin 78.43 52.85 46.21 47.76 89.33
DISCO RS Med AD Sci Fin
RS 78.57
Med 75.80 52.36
AD 76.37 49.78 53.04
Sci 76.11 45.76 49.26 49.23
Fin 76.03 45.20 43.79 42.33 88.95
DISCO RS Med AD Sci Fin
RS 79.20
Med 77.96 52.44
AD 78.05 49.85 53.03
Sci 77.26 46.32 53.08 51.99
Fin 77.78 46.25 50.45 49.51 89.71
C.2 Baseline Results in Ability Continual Learning
Table 15: Result matrices of LLaVA-based baselines in ability continual learning. denotes the original method with replay data.
LoRA-FT OCR Math VP APP
OCR 33.30
Math 32.60 34.20
VP 31.70 32.80 65.10
APP 23.60 33.70 55.84 32.50
LoRA-FT OCR Math VP APP
OCR 32.60
Math 33.60 33.80
VP 31.10 33.50 66.12
APP 21.80 32.70 58.38 28.75
O-LoRA OCR Math VP APP
OCR 32.90
Math 29.80 33.60
VP 27.40 33.70 58.63
APP 29.60 32.90 52.41 33.75
O-LoRA OCR Math VP APP
OCR 34.00
Math 28.40 36.80
VP 28.90 33.90 61.55
APP 29.60 31.30 60.79 27.50
MoELoRA OCR Math VP APP
OCR 32.70
Math 32.50 33.30
VP 30.80 33.00 64.59
APP 26.70 32.80 56.85 27.22
MoELoRA OCR Math VP APP
OCR 32.70
Math 29.40 33.10
VP 32.60 32.50 65.61
APP 19.80 32.20 54.19 30.00
CL-MoE OCR Math VP APP
OCR 33.00
Math 32.30 33.60
VP 30.20 32.50 64.72
APP 19.90 32.70 53.43 30.69
CL-MoE OCR Math VP APP
OCR 33.20
Math 34.30 36.70
VP 32.00 33.20 64.97
APP 25.40 31.80 60.91 30.00
HiDe OCR Math VP APP
OCR 33.40
Math 30.90 32.80
VP 30.40 33.30 56.98
APP 24.60 32.10 46.32 28.75
HiDe OCR Math VP APP
OCR 34.10
Math 32.60 35.70
VP 30.70 32.60 53.81
APP 24.60 28.40 30.71 23.75
SEFE OCR Math VP APP
OCR 33.00
Math 32.20 32.60
VP 31.80 33.30 64.59
APP 26.00 33.40 57.74 33.75
SEFE OCR Math VP APP
OCR 33.60
Math 33.80 37.50
VP 32.80 36.10 66.50
APP 25.60 34.80 57.61 31.39
DISCO OCR Math VP APP
OCR 32.90
Math 31.80 33.40
VP 31.00 34.50 59.64
APP 32.90 33.10 60.15 30.14
DISCO OCR Math VP APP
OCR 33.40
Math 32.10 36.60
VP 32.20 37.00 63.07
APP 34.20 35.00 61.55 27.50
Table 16: Result matrices of InternVL-based baselines in ability continual learning. denotes the original method with replay data.
LoRA-FT OCR Math VP APP
OCR 32.20
Math 33.10 33.30
VP 31.80 32.30 68.02
APP 21.40 32.80 60.28 29.86
LoRA-FT OCR Math VP APP
OCR 31.60
Math 35.30 35.40
VP 32.60 31.10 68.27
APP 26.30 34.20 62.56 31.25
O-LoRA OCR Math VP APP
OCR 32.70
Math 31.10 34.20
VP 30.20 33.00 63.20
APP 25.50 32.30 64.59 24.44
O-LoRA OCR Math VP APP
OCR 34.00
Math 30.90 34.40
VP 31.00 33.20 65.86
APP 21.70 31.10 59.77 31.25
MoELoRA OCR Math VP APP
OCR 32.20
Math 29.90 33.30
VP 29.20 32.80 67.64
APP 17.20 32.70 55.33 32.50
MoELoRA OCR Math VP APP
OCR 32.90
Math 31.50 36.50
VP 30.90 32.30 65.74
APP 13.90 29.70 54.95 32.50
HiDe OCR Math VP APP
OCR 33.40
Math 26.30 33.60
VP 30.10 33.00 61.80
APP 17.70 33.00 41.12 20.28
HiDe OCR Math VP APP
OCR 33.40
Math 28.10 34.70
VP 31.10 32.20 55.33
APP 25.30 29.20 42.13 20.28
DISCO OCR Math VP APP
OCR 31.90
Math 31.70 34.00
VP 32.10 33.50 63.45
APP 30.60 33.10 65.36 27.50
DISCO OCR Math VP APP
OCR 34.70
Math 31.50 34.70
VP 31.50 34.60 62.31
APP 32.30 32.30 64.97 30.14
C.3 Detailed Results of MR-LoRA
Table 17: Result matrices of MR-LoRA in domain continual learning. LLaVA denotes LLaVA-based MR-LoRA, and InternVL denotes InternVL-based MR-LoRA.
LLaVA RS Med AD Sci Fin
RS 81.06
Med 81.06 65.73
AD 81.06 65.71 54.17
Sci 81.06 65.68 54.17 56.11
Fin 80.87 65.32 54.12 56.71 91.12
InternVL RS Med AD Sci Fin
RS 81.49
Med 81.49 66.40
AD 81.49 66.42 54.56
Sci 81.47 65.81 54.56 54.05
Fin 81.48 65.80 54.56 54.40 91.07
Table 18: Result matrices of MR-LoRA in ability continual learning. LLaVA denotes LLaVA-based MR-LoRA, and InternVL denotes InternVL-based MR-LoRA.
LLaVA OCR Math VP APP
OCR 33.60
Math 33.50 36.50
VP 33.50 36.40 64.97
APP 33.70 36.20 65.10 32.50
InternVL OCR Math VP APP
OCR 32.20
Math 33.80 36.40
VP 33.30 35.60 67.89
APP 33.00 35.70 67.51 33.75
Appendix D Limitations and Broader Impacts
D.1 Limitations
Although our study makes valuable contributions, we acknowledge the following limitations: (1) Model size and training limitations: This research focuses exclusively on MLLMs with 7 billion parameters. Owing to computational constraints, we did not explore larger models. (2) potential inaccuracies in the StockQA dataset: Our StockQA dataset is generated by Qwen2.5-VL (Bai et al., 2025), and the model may inadvertently produce inaccurate or misleading data. Moreover, biases inherent in the training data could manifest in the generated dataset, influencing the outcomes and interpretations of subsequent analyses. We hope to address these limitations in our future work to build a practical and lifelong-evolving MLLM.
D.2 Broader Impacts
Positively, such work advances the ability of AI systems to learn adaptively from ongoing streams of diverse data, enabling applications in education, assistive technologies, and personalized healthcare. These systems could provide more context-aware and accessible tools that evolve over time to better support users’ needs. Moreover, robust continual learning reduces the need for retraining from scratch, leading to more energy-efficient and sustainable AI development. However, there are potential negative impacts. Without careful design, continual learning systems may inadvertently retain or amplify biases from evolving data streams, leading to fairness concerns. The dynamic nature of these models also complicates auditing and accountability, as their behavior changes over time. Additionally, if misused, adaptive models could enhance surveillance or manipulation by continuously tailoring outputs to influence user behavior. To mitigate these risks, transparency, rigorous evaluation, and ethical safeguards must be integrated into both benchmark design and method development.
Appendix E Inference Optimization with Caching
A key advantage of our method is its computational efficiency during inference. While our approach involves two distinct phases, we introduce a caching strategy that collapses the computational overhead. The most intensive operation—the forward pass through the backbone network (i.e., the visual encoder and LLM) is performed only once. We cache the resulting hidden states from each layer (specifically, the KV cache) after this single pass. Subsequently, our two lightweight modules, the router and the expert LoRA, operate sequentially on these cached states, obviating the need for a second full forward pass. This optimization reduces the computational cost from that of two full inferences to only marginally more than a single one, achieving a practical deployment cost comparable to standard single-pass methods, such as LoRA-FT (Hu et al., 2021).
Appendix F Use of LLM
In the preparation of this manuscript, we utilized a Large Language Model (LLM) in a capacity analogous to a conventional grammar-checking tool. Its application was strictly confined to copy-editing tasks, such as correcting spelling, improving grammar, and enhancing the clarity of author-generated text. No part of the research ideation, methodology, data analysis, or generation of substantive content was performed by the LLM.
Appendix G Visualization
G.1 Illustration of MLLM-CL Benchmark
In this section, we show more examples of our MLLM-CL benchmark in domain continual learning and ability continual learning.
Figure 15: Examples of remote sensing task in domain continual learning.
Figure 16: Examples of medical task in domain continual learning.
Figure 17: Examples of science task in domain continual learning.
Figure 18: Examples of autonomous driving task in domain continual learning.
Figure 19: Examples of OCR task in ability continual learning.
Figure 20: Examples of math task in ability continual learning.
Figure 21: Examples of GUI agent task in ability continual learning.
Figure 22: Examples of visual perception task in ability continual learning.
G.2 Visualization of Results
Fig.˜23 provides examples during DCL and ACL, respectively. We can find that some baselines like LoRA (Hu et al., 2021), MoELoRA (Chen et al., 2024a), HiDE (Guo et al., 2025a) overfit to the last learned task and output options that do not exist in domain continual learning. In ACL, most baselines, including HiDe (Guo et al., 2025a), DISCO (Guo et al., 2025b), CL-MoE (Huai et al., 2025), etc., miss part of their OCR ability and do not answer the question correctly.
Figure 23: Visualization of MR-LoRA and other baselines under domain continual learning and ability continual learning. The left part is testing the autonomous driving task after learning all domain tasks, while the right part is testing the OCR tasks after learning all ability tasks.